OpenStudy (nudeoua):

A rectangle has a length of the fifth root of 16 inches and a width of 2 to the 1 over 5 power inches. Find the area of the rectangle. 2 to the 3 over 5 power inches squared 2 to the 4 over 5 power inches squared 2 inches squared 2 to the 2 over 5 power inches squared

OpenStudy (er.mohd.amir):

area = l*b= =16^(1/5)*2^(1/5) =ans use 16=2^4 16^(1/5)=2^(4/5)

OpenStudy (fouzberzerk):

Hey, AMIR I could use some help on a new question.

OpenStudy (fouzberzerk):

I posted it, it's the top question.

Directrix (directrix):

@nudeoua The correct answer is 2.